OpenStudy (junyang96):

Probability. If X and Y are events of a trial with P(X)=1/2, P(X|Y')=2/3, P(X|Y)=3/7 find P(Y)

OpenStudy (junyang96):

Got it. P(Y)=7/10

ganeshie8 (ganeshie8):

how ?

OpenStudy (junyang96):

Since \[P(X|Y')=\frac{ 2 }{ 3 }=\frac{ P(XintersectY') }{ P(Y') }\] so P(Y')= 3/total \[P(X|Y)=\frac{ 3 }{ 7 }=\frac{ P(XintersectY) }{ P(Y) }\] P(Y)=7/total We know P(Y)+P(Y')=1 7/total +3/total= 1 total =7+3=10 so P(Y)=0.7

OpenStudy (junyang96):

I'd change the "total" to "sample space" to avoid confusion